# Improving Autoformalization Using Direct Dependency Retrieval

### v1

@@ Abstract
The convergence of deep learning and formal mathematics has spurred research in formal verification. Statement autoformalization, a crucial first step in this process, aims to translate informal descriptions into machine-verifiable representations but remains a significant challenge. The core difficulty lies in the fact that existing methods often suffer from a lack of contextual awareness, leading to hallucination of formal definitions and theorems. Furthermore, current retrieval-augmented approaches exhibit poor precision and recall for formal library dependency retrieval, and lack the scalability to effectively leverage ever-growing public datasets. To bridge this gap, we propose a novel retrieval-augmented framework based on DDR (Direct Dependency Retrieval) for statement autoformalization. Our DDR method directly generates candidate library dependencies from natural language mathematical descriptions and subsequently verifies their existence within the formal library via an efficient suffix array check. Leveraging this efficient search mechanism, we constructed a dependency retrieval dataset of over 500,000 samples and fine-tuned a high-precision DDR model. Experimental results demonstrate that our DDR model significantly outperforms SOTA methods in both retrieval precision and recall. Consequently, an autoformalizer equipped with DDR shows consistent performance advantages in both single-attempt accuracy and multi-attempt stability compared to models using traditional selection-based RAG methods.
Machine Learning, ICML
@@ 1 Introduction
Theorem provers, such as Lean (moura2021lean), Coq (bertot2013interactive), and Isabelle (nipkow2002isabelle), are symbolic systems designed for formal verification whose soundness and completeness are proven in theory. Grounded in specific foundational logics, these systems employ proprietary formal languages to encode mathematical concepts, statements, and proofs. The logical kernel within each system can then verify the validity of these formalized objects with complete rigor (nawaz2019survey). However, this process eschews the flexibility of natural language, demanding that users not only master the strict formal syntax but also possess a profound understanding of the semantic mapping between mathematical ideas and their formal representations. Consequently, the effective utilization of these theorem provers presents a formidable challenge to researchers (li2024survey).
@@
Statement autoformalization refers to the process of automatically translating mathematical statements from natural language into a formal language comprehensible to theorem provers (liu2025rethinking). As a foundational step in the automated theorem proving, the correctness of its output is paramount. It determines whether the proof target of a subsequent prover is precisely aligned with the intent of the original statement. Beyond this core function, statement autoformalization exhibits significant potential for broader applications and innovation. These applications include synthesizing training data for neural theorem provers (lin2025goedel), automated program verification (lin2024fvel), curating and filtering informal reasoning datasets (zhang2025deeptheorem), and evaluating and enhancing the informal reasoning capabilities of large language models (e.g., through techniques like rejection sampling (zhou2024don)).
@@
Current mainstream research in autoformalization primarily encompasses three direction. First, constructing larger and more challenging datasets for SFT (Supervised Fine-Tuning) (cao2025informal; peng2025criticlean; yu2025formalmath; miranda2025veribench; kimina_prover_2025; ying2024lean; zheng2021minif2f; azerbayev2023proofnet); second, encoding and indexing the libraries of theorem provers to support retrieval (liu2025rethinking; yang2023leandojo; wang2025aria); and third, employing methods like ICL (In-Context Learning) to enable LLMs (Large Language Models) to directly translate informal statements into formal code (wu2022autoformalization; zhang2024consistent).
@@
However, these approaches face two critical limitations. First, relying solely on generative models to directly synthesize formal code often overlooks the context of existing definitions and lemmas within the formal library (yang2023leandojo; zhang2024consistent). This lack of contextual awareness is a primary reason for hallucinations in the model’s use of identifiers and construction of syntax . For complex theorems, effectively leveraging relevant conclusions from the library is crucial for reducing formalization difficulty. Without this capability, LLMs are forced to locally reinvent functionally equivalent but more complex expressions or to selectively ignore key premises they cannot handle, which readily introduces errors. Similiar results have been reported in wu2022autoformalization and yu2025formalmath, pointing out that definitional mismatches between informal mathematics and the formal library, as well as the inappropriate usage of tactics and theorems are primary causes of formalization failures. Figure 1 (a) intuitively illustrates this issue: unable to access the predefined Nat.factorization in the library, the model attempts to construct an equivalent operator locally, resulting in erroneous code. Similarly, the results in Table 3, which show a significant performance decline for all ICL-based models as task difficulty increases, further underscore the severity of this problem. Second, the retrieval strategies for library dependencies in existing methods are often rudimentary, yielding results with weak semantic relevance to the informal statement (azerbayev2023proofnet; yang2023leandojo). The libraries of mainstream theorem provers are vast and continually growing; for instance, Lean 4’s mathlib contains over 240,000 entries, including many with similar descriptions and usage. Simple retrieval methods struggle to achieve both high efficiency and precision. This challenge is confirmed by the experiments in liu2025rethinking as well as our own results presented in Table 2.
@@
To address the aforementioned challenges, we propose a novel dependency retrieval method DDR, and the corresponding framework for statement autoformalization. In this context, dependency retrieval aims to identify relevant formal objects, such as definitions and theorems, from a theorem prover’s library based on an informal mathematical statement. Unlike traditional selection-based retrieval methods, DDR leverages the contextual understanding of LLMs to directly generate potential formal dependencies. To empower the LLM with this capability, we have designed a suffix array-based methodology for both data generation and dependency verification. Specifically, this process begins by constructing an efficient database from the theorem prover’s library using a suffix array. Subsequently, leveraging existing informal-formal statement pairs from public datasets, we extract candidate dependencies from the formal code and verify them against the database via the efficient binary search. This produces high-quality labels for fine-tuning the DDR model. The same verification mechanism is reused at inference time to validate the dependencies generated by the DDR model. Given the ever-increasing size of modern theorem prover libraries, we place special emphasis on the efficiency of the verification step in our RAG pipeline. Let $N$ denote the volume of the library, the time complexities for database construction and verification in our proposed method are $\mathcal{O}(N)$ and $\mathcal{O}(\log N)$ respectively, ensuring the scalability of our approach to libraries of any size. By applying the above process to the FineLeanCorpus dataset, we successfully extracted all Mathlib dependencies for over 500,000 samples, reducing the processing time from approximately 2.5 days to under 2 hours. This achievement has enabled the construction of a large-scale dataset for fine-tuning and paves the way for the online deployment of RAG services. Experimental results demonstrate that DDR significantly improves both recall and precision in dependency retrieval compared to traditional methods. Consequently, an autoformalizer equipped with DDR exhibits consistent advantages in both single-attempt accuracy and multi-attempt stability over autoformalizers using conventional RAG approaches.
@@
    We propose a library dependency verification process based on suffix array. With its $\mathcal{O}(\log N)$ verification time complexity, this highly efficient and scalable method provides critical support for deploying online RAG services and constructing large-scale fine-tuning datasets.
@@
    We introduce an innovative RAG method based on a generation-plus-verification paradigm. The core component, DDR (Direct Dependency Retrieval), unleashes the contextual awareness of LLMs, enabling them to directly generate potential library dependencies from informal input, enhancing the precision and recall of retrieval results.
@@
    We conducted comprehensive experiments on statement autoformalization. The results validate that DDR consistently outperforms baseline models that rely on traditional retrieval methods.
@@
    We have constructed and open-sourced a library dependency dataset with over 500,000 samples. Each sample consists of an informal statement and its corresponding formal library dependencies, suitable for fine-tuning LLMs. The methodology for constructing this dataset is generalizable and can be applied to other formal corpora, fostering more diverse research in statement autoformalization.
Figure 1: Comparison of two paradigms for formal dependency retrieval. (a) Select-based retrieval using embedding cosine similarity: both the input statement and library items are encoded into embeddings, and the top-k items with highest cosine similarity are retrieved as dependencies. (b) DDR (Direct Dependency Retrieval): an LLM generates potential dependencies based on context, followed by a suffix-array-based verification step. DDR achieves higher retrieval accuracy and scalability with a more direct and efficient workflow.
@@ 2 Related Work
Research on evaluating the fidelity of autoformalization has explored various methods, such as manual expert review (wang2018first; peng2025criticlean), LLM review (yang2023leandojo; sidhu2024evaluation), auxiliary metrics like Perplexity (wang2018first) and BLEU (papineni2002bleu; wang2018first), and type checking (poiroux2024improving) within a theorem prover. However, these approaches have inherent limitations in capturing the deep semantic equivalence between informal and formal mathematical statements. They may overlook semantic inconsistencies between the two modalities or produce false negatives by incorrectly rejecting semantically valid formalizations (liu2025rethinking). To overcome this, our work adopts BEq (Bidirectional Extender Definitional Equivalence) as the primary evaluation metric (liu2025rethinking). This choice is motivated by BEq’s high alignment with human expert judgment in assessing semantic equivalence. The core principle of BEq is to define two formal statements as equivalent if they are mutually transformable through a predefined, finite, and verifiable set of transformation rules (aczel2013homotopy).
Figure 2: Workflow of the suffix array auxiliary checking process. All library items are concatenated into a single string with delimiters, followed by suffix array construction. Each pending item is then queried via binary search to determine its match status (exact, partial, or none). This mechanism supports both dataset construction and LLM output verification.
@@ 3.1 Direct Dependency Retrieval
To prevent hallucination during generation, existing dependency retrieval methods select items from a pre-existing library to form their output (yang2023leandojo; liu2025rethinking), as illustrated in Figure 1 (a). However, this selected-based paradigm suffers from two primary limitations. First, training an encoder to produce effective embeddings for every formal object in the library to maximize retrieval performance is a challenge in itself (xiong2020approximate; gao2023retrieval). Second, the method is sensitive to superficial variations in the input: semantically equivalent mathematical statements may yield different results in maximum cosine similarity retrieval due to minor differences in their syntactic structure or notational representation (asl2024robustsentembed). While the first issue can be partially mitigated by techniques such as optimizing the input format for the embedding model, the second remains an challenge. Our empirical results confirm that such embedding-based retrieval methods exhibit low recall and precision. Inaccurate retrieval results don’t contribute to the language model in autoformalization tasks. Figure 1 (a) illustrates this failure mode: in the absence of a usable dependency from the retrieved Mathlib candidates, the model attempts to construct a local function functionally equivalent to factorization, leading to an erroneous output. Furthermore, this approach lacks scalability, as it cannot effectively leverage the growing volume of available datasets to continuously enhance model capabilities.
@@
To address the challenge of LLMs hallucinating non-existent or irrelevant library dependencies during autoformalization, we propose a generation-plus-verification pipeline, with DDR as its core component. DDR is a LLM-based SFT model to operate in an end-to-end fashion, directly extracting potential library dependencies from an input mathematical statement. In contrast to conventional two-stage retrieval methods that require additional encoder model (liu2025rethinking), the design of DDR streamlines the architecture, leading to a simpler, more efficient, and better-performing approach. Our experimental results demonstrate the effectiveness of DDR, showing significant improvements in both recall and precision compared to mainstream embedding-based retrieval methods. Accurate dependency retrieval reduces the complexity of the subsequent formalization task, thereby enhancing the overall performance of the autoformalizer. As illustrated in Figure 1 (b), providing the model with precise dependencies, such as Nat.factorization and Nat.factorial, enables it to more reliably generate the correct formal output, thus increasing the overall success rate. Furthermore, the DDR framework is designed to leverage existing datasets to augment its training corpus, providing a viable path for the continuous improvement of its retrieval accuracy.
The effective application of the DDR method confronts two primary challenges. First, unlike embedding-based approaches that retrieve existing entries from a library, DDR relies on a LLM to directly generate potential dependencies. This generative nature necessitates the verification of its output to filter out non-existent items not present in the formal library (i.e., hallucinations). The complexity of this verification process is exacerbated by name ambiguity: as the model is trained on code, its output often consists of unqualified short names (e.g., sqrt) rather than the fully qualified names of formal objects (e.g., Int.sqrt). Such short names can correspond to multiple entities within the library (e.g., Nat.sqrt or Real.sqrt), posing a challenge for precise matching. Second, training a high-performance DDR model requires specialized datasets. However, existing corpora of formalized mathematics generally lack high-quality paired data that directly links mathematical statements to their required dependencies. Consequently, how to efficiently construct such training data from existing formal code and effectively leverage the ever-growing new corpora presents a critical problem. To systematically address these issues of verification and data construction, we will elaborate on our proposed solutions in the subsequent two subsections.
3.2 Suffix Array Auxiliary

### v2

@@ Abstract
Statement autoformalization, a crucial first step in formal verification, aims to transform informal descriptions of math problems into machine-verifiable formal representations but remains a significant challenge. The core difficulty lies in the fact that existing language models often suffer from a lack of contextual awareness, leading to hallucination of formal dependencies such as definitions and theorems. Current dependency retrieval approaches exhibit poor precision and recall, and lack the scalability to leverage ever-growing public datasets. To bridge this gap, we propose a novel retrieval-augmented framework based on Direct Dependency Retrieval (DDR). DDR directly generates candidate formal dependencies from natural-language mathematical descriptions and verifies their existence in the formal library via an efficient Suffix Array Check (SAC). Built on a SAC-constructed dependency retrieval dataset of over 500,000 samples, a high-precision DDR model is fine-tuned and shown to significantly outperform state-of-the-art methods in both retrieval precision and recall, leading to superior advantage in the autoformalization tasks. SAC also contributes in assessing formalization difficulty and enabling explicit quantification of the hallucination in In-Context Learning (ICL).
Improving Autoformalization Using Direct Dependency Retrieval
Shaoqi Wang   Lu Yu   Siwei Lou   Feng Yan   Chunjie Yang Zhejiang University, {sq_w,swlou,cjyang999}@zju.edu.cn Ant Group Hong Kong Polytechnic University, {feng555.yan}@polyu.edu.hk
@@ 1 Introduction
Theorem provers, such as Lean (moura2021lean), Coq (bertot2013interactive), and Isabelle (nipkow2002isabelle), are symbolic systems designed for formal verification whose soundness and completeness are proven. Grounded in specific foundational logics, these systems employ proprietary formal languages to encode mathematical statements, definitions, and proofs. The logical kernel within each system can then verify the validity of these formalized objects with complete rigor (nawaz2019survey). However, this process eschews the flexibility of natural language, demanding that users not only master the strict formal syntax but also possess a profound understanding of the semantic mapping between mathematical ideas and their formal representations. Consequently, the effective utilization of these theorem provers presents a formidable challenge to researchers (li2024survey).
@@
Statement autoformalization refers to the process of automatically translating mathematical statements from natural language into a formal language comprehensible to theorem provers (liu2025rethinking). As a foundational step in the formal verification, the correctness of its output is paramount. It determines whether the proof target is precisely aligned with the intent of the original statement. Beyond this core function, statement autoformalization exhibits significant potential for broader applications and innovation. These applications include synthesizing training data for neural theorem provers (lin2025goedel), automated program verification (lin2024fvel), and curating and filtering informal reasoning datasets (zhang2025deeptheorem). , and evaluating and enhancing the informal reasoning capabilities of LLMs (e.g., through techniques like rejection sampling (zhou2024don)).
@@
Current mainstream research in autoformalization primarily encompasses three direction. First, constructing larger and more challenging datasets for Supervised Fine-Tuning (SFT) (cao2025informal; peng2025criticlean; yu2025formalmath; miranda2025veribench; kimina_prover_2025; ying2024lean; zheng2021minif2f; azerbayev2023proofnet); second, encoding and indexing the libraries of theorem provers to support retrieval (liu2025rethinking; yang2023leandojo; wang2025aria); and third, employing methods like ICL to enable Large Language Models (LLMs) to directly translate informal statements into formal code (wu2022autoformalization; zhang2024consistent).
Figure 1: Comparison of two paradigms for formal dependency retrieval. (a) Select-based retrieval using embedding cosine similarity: both the input statement and library items are encoded into embeddings, and the top-k items with highest cosine similarity are retrieved as dependencies. (b) DDR generates potential dependencies based on context, followed by SAC verification step.
@@
However, these approaches face two critical limitations. First, relying solely on generative models to directly synthesize formal code often overlooks the existing definitions and lemmas within the formal library (yang2023leandojo; zhang2024consistent). This ignorance is a primary reason for the model to hallucinate non-existent formal objects or generate syntactically incorrect code (wu2022autoformalization; liu2025rethinking). Results presented in Table 2 demonstrate that even with a two-stage ICL approach, where dependencies are retrieved prior to the main task, the problem of hallucination in model outputs remains significant. Second, the retrieval strategies for library dependencies in existing methods are rudimentary, yielding results with weak semantic relevance to the informal statement (azerbayev2023proofnet; yang2023leandojo). The libraries of mainstream theorem provers are vast and continually growing; for instance, Lean 4’s mathlib contains over 240,000 entries, including many with similar descriptions and usage. Simple retrieval methods struggle to achieve both high efficiency and precision. This challenge is confirmed by the experiments in liu2025rethinking as well as our own results presented in Table 3.
@@
To address the aforementioned challenges, we propose a novel dependency retrieval method DDR, and the corresponding framework for statement autoformalization. In this context, dependency retrieval aims to identify relevant formal objects, such as definitions and theorems, from a theorem prover’s library based on an informal mathematical statement (liu2025rethinking). DDR leverages the contextual understanding of LLMs to directly generate potential formal dependencies. To empower the LLM with this capability, we have designed SAC for both data generation and dependency verification. Specifically, this process begins by constructing an efficient database from the theorem prover’s library using a suffix array. Subsequently, leveraging existing informal-formal statement pairs from public datasets, we extract candidate dependencies from the formal code and verify them against the database via the efficient binary search. This produces high-quality labels for fine-tuning the DDR model. The same verification mechanism is reused at inference time to validate the dependencies generated by the DDR model. Experimental results demonstrate that DDR significantly improves both recall and precision in dependency retrieval, leading to superior advantage in the autoformalization tasks.
@@
    We propose DDR as a novel generation-plus-verification dependency retrieval paradigm in autoformalization tasks, and SAC as a flexible dependency verification method.
@@
    We provide a quantitative analysis of ICL for dependency retrieval, revealing its inherent and severe hallucination problems.
@@
    We conduct comprehensive experiments on statement autoformalization. The results validate that DDR can serve as a plug-and-play module to enhance the performance of existing autoformalizers and ICL methods.
@@
    The proposed SAC provides a potential sound mechanism for assessing formalization difficulty, overcomes the inherent subjectivity and stochasticity of LLM-based graders.
  5. 5.
    We have constructed and open-sourced a library dependency dataset with over 500,000 samples. The methodology for constructing this dataset is generalizable and can be applied to other formal corpora, fostering more diverse research in statement autoformalization.
@@ 2 Related Work
Figure 2: Schematic diagram of SAC process. All library items are concatenated into a single string with delimiters, followed by suffix array construction. Each pending item is then queried via binary search to determine its match status (exact, partial, or none). This mechanism supports both dataset construction and DDR output verification.
@@
Research on evaluating the fidelity of autoformalization has explored various methods, such as manual expert review (wang2018first; peng2025criticlean), LLM review (yang2023leandojo; sidhu2024evaluation), auxiliary metrics like Perplexity (wang2018first), BLEU (papineni2002bleu; wang2018first), and type checking (poiroux2024improving) within a theorem prover. However, these approaches have inherent limitations in capturing the deep semantic equivalence between informal and formal mathematical statements. They may overlook semantic inconsistencies between the two modalities or produce false negatives by incorrectly rejecting semantically valid formalizations (liu2025rethinking). To overcome this, our work adopts Bidirectional Extender Definitional Equivalence (BEq) as the primary evaluation metric (liu2025rethinking). This choice is motivated by BEq’s high alignment with human expert judgment in assessing semantic equivalence. The core principle of BEq is to define two formal statements as equivalent if they are mutually transformable through a predefined, finite, and verifiable set of transformation rules (aczel2013homotopy).
@@ 3.1 Direct Dependency Retrieval
To prevent hallucination during generation, existing dependency retrieval methods select items from a pre-existing library to form their output (yang2023leandojo; liu2025rethinking), as illustrated in Figure 1 (a). However, this selected-based paradigm suffers from two primary limitations. First, training an encoder to produce effective embeddings for every formal object in the library to maximize retrieval performance is a challenge in itself (xiong2020approximate; gao2023retrieval). Second, the method is sensitive to superficial variations in the input: semantically equivalent mathematical statements may yield different results in maximum cosine similarity retrieval due to minor differences in their syntactic structure or notational representation (asl2024robustsentembed). Our empirical results in Table 3 confirm that such embedding-based retrieval methods exhibit low recall and precision. Inaccurate retrieval results don’t contribute to the language model in autoformalization tasks. Figure 1 (a) illustrates this failure mode: in the absence of a usable dependency from the retrieved Mathlib candidates, the model attempts to construct a local function functionally equivalent to factorization, leading to an erroneous output. Furthermore, this approach lacks scalability, as it cannot effectively leverage the growing volume of available datasets to continuously enhance model capabilities.
Method    Diff01    Diff23    Diff45    Diff67    Diff89 $_{\text{m}} Hall$$_{\text{std}}$ $_{\text{m}} Hall$$_{\text{std}}$ $_{\text{m}} Hall$$_{\text{std}}$ $_{\text{m}} Hall$$_{\text{std}}$ $_{\text{m}} Hall$$_{\text{std}}$ Claude-3.5-Sonnet 0.349 842 767 296 226 50.349\,842\,767\,296\,226\,5 0.321 705 896 805 355 60.321\,705\,896\,805\,355\,6 0.300 598 443 2230.300\,598\,443\,223 0.257 305 310 183 968 60.257\,305\,310\,183\,968\,6 0.277 148 809 5230.277\,148\,809\,523 0.258 708 434 626 845 840.258\,708\,434\,626\,845\,84 0.284 793 650 793 500 050.284\,793\,650\,793\,500\,05 0.220 764 928 185 668 670.220\,764\,928\,185\,668\,67 0.276 160 173 1630.276\,160\,173\,163 0.221 181 722 747 262 330.221\,181\,722\,747\,262\,33 DeepSeek-R1 0.234 589 265 257 731 940.234\,589\,265\,257\,731\,94 0.282 361 595 375 512 250.282\,361\,595\,375\,512\,25 0.330 162 841 356 923 10.330\,162\,841\,356\,923\,1 0.313 336 322 060 300 80.313\,336\,322\,060\,300\,8 0.265 354 183 465 656 60.265\,354\,183\,465\,656\,6 0.289 163 096 881 404 20.289\,163\,096\,881\,404\,2 0.306 544 110 086 50.306\,544\,110\,086\,5 0.297 199 460 504 328 20.297\,199\,460\,504\,328\,2 0.302 954 725 8290.302\,954\,725\,829 0.288 580 908 324 351 80.288\,580\,908\,324\,351\,8 GPT-4o 0.290 530 303 031 818 20.290\,530\,303\,031\,818\,2 0.279 890 001 371 935 650.279\,890\,001\,371\,935\,65 0.298 034 188 031 282 070.298\,034\,188\,031\,282\,07 0.274 865 951 589 759 10.274\,865\,951\,589\,759\,1 0.287 544 867 191 959 80.287\,544\,867\,191\,959\,8 0.265 736 723 207 650 960.265\,736\,723\,207\,650\,96 0.290 236 898 778 392 040.290\,236\,898\,778\,392\,04 0.278 007 943 706 211 30.278\,007\,943\,706\,211\,3 0.300 733 890 125 757 630.300\,733\,890\,125\,757\,63 0.258 576 099 806 656 40.258\,576\,099\,806\,656\,4 Ling-flash-2.0 0.326 612 785 787 1560.326\,612\,785\,787\,156 0.316 176 237 531 755 40.316\,176\,237\,531\,755\,4 0.385 175 652 200 500 040.385\,175\,652\,200\,500\,04 0.314 393 792 807 857 730.314\,393\,792\,807\,857\,73 0.342 384 920 637 000 10.342\,384\,920\,637\,000\,1 0.298 188 535 254 873 270.298\,188\,535\,254\,873\,27 0.325 408 019 2680.325\,408\,019\,268 0.276 772 886 385 413 10.276\,772\,886\,385\,413\,1 0.362 064 913 029 000 070.362\,064\,913\,029\,000\,07 0.276 638 773 854 510 90.276\,638\,773\,854\,510\,9 Qwen-Max 0.229 807 692 306 730 760.229\,807\,692\,306\,730\,76 0.302 868 995 190 336 670.302\,868\,995\,190\,336\,67 0.274 352 489 913 265 330.274\,352\,489\,913\,265\,33 0.305 769 901 832 063 040.305\,769\,901\,832\,063\,04 0.209 369 488 533 333 340.209\,369\,488\,533\,333\,34 0.241 521 686 559 3190.241\,521\,686\,559\,319 0.241 563 880 005 025 130.241\,563\,880\,005\,025\,13 0.264 552 600 591 864 860.264\,552\,600\,591\,864\,86 0.247 436 120 550 753 790.247\,436\,120\,550\,753\,79 0.238 268 638 411 097 50.238\,268\,638\,411\,097\,5 R@5 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 R@10 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 DDR $0.016$ 0.083 891 822 234 576 820.083\,891\,822\,234\,576\,82 0.006 027 253 668 553 460.006\,027\,253\,668\,553\,46 0.036 863 586 454 198 930.036\,863\,586\,454\,198\,93 0.004 540 420 820 155 0380.004\,540\,420\,820\,155\,038 0.026 243 591 065 426 320.026\,243\,591\,065\,426\,32 0.003 062 678 062 820 512 80.003\,062\,678\,062\,820\,512\,8 0.022 527 428 053 560 6230.022\,527\,428\,053\,560\,623 0.013 735 240 031 097 560.013\,735\,240\,031\,097\,56 0.087 415 728 148 786 620.087\,415\,728\,148\,786\,62
[…]
@@
To address the challenge of LLMs hallucinating non-existent or irrelevant library dependencies during autoformalization, we propose a generation-plus-verification pipeline, with DDR as its core component. DDR is a SFT model to operate in an end-to-end fashion, directly extracting potential library dependencies from an input mathematical statement. The SFT based paradigm enables it to effectively absorb and leverage continuously growing training data, thereby demonstrating excellent scalability. In contrast to conventional two-stage retrieval methods that require additional encoder model (liu2025rethinking), the design of DDR streamlines the architecture, leading to a simpler, more efficient, and better-performing approach. Accurate dependency retrieval reduces the complexity of the subsequent formalization task, thereby enhancing the overall performance of the autoformalizer. As illustrated in Figure 1 (b), providing the model with precise dependencies enables it to more reliably generate the correct formal output, thus increasing the overall success rate.
The effective application of the DDR method entails two primary challenges: first, constructing a high-quality dataset for model fine-tuning, and second, implementing an output filtering mechanism to validate the generated dependencies. Such challenges can be addressed by the following SAC method.
3.2 Suffix Array Check